OpenStudy (anonymous):

Can you help me real quick? Choose the slope-intercept equation of the line that passes through the point (4,-6) and is parallel to y=-1/2x+5. A.y=-2x+2 B.y=2x-14 C.y=1/2x-8 D.y=-1/2x-4

OpenStudy (anonymous):

First of all note the point A(4,-6) then use the "point-slope formula"; which bring to: y-yA=m(x-xA)=>y+6=m(x-4)=>mx-y-4m-6=0; if the two lines are parallel then their slopes must be equal so m=-1/2 =>(-1/2)x-y+2-6=0=> ->(-1/2)x-y-4=0<-

OpenStudy (anonymous):

So the final answer is D.y=-1/2x-4